OpenStudy (anonymous):

(a4-b4) divided by (a2-b2) polynomial question. Plz get it solved & kindly reply.

OpenStudy (lalaly):

\[a^4-b^4=(a^2)^2-(b^2)^2=(a^2-b^2)(a^2+b^2)\]

Parth (parthkohli):

But this can be simplified further @lalaly \(\Large \color{MidnightBlue}{\Rightarrow (a + b)(a - b)(a^2 + b^2) }\)

OpenStudy (lalaly):

yeah but u dont have to since its divided by (a^2-b^2) "check the question"

Parth (parthkohli):

Oh yes.

OpenStudy (anonymous):

But the book answer is only (a^2+b^2) this is for lalaly to plz re-check .

OpenStudy (lalaly):

yeah i didnt give u the last answer,, i just showd u another way of writing the numerator so it can be easier for u to figure it out... YOU recheck my post

OpenStudy (lalaly):

i didnt divide yet ,, when u divide, (a^2-b^2) on the top cancels with (a^2-b^2) in the bottom ,,, so u are left with (a^2+b^2)

OpenStudy (anonymous):

Thanks for your kind collaboration

OpenStudy (anonymous):

7 puto http://t.qkme.me/3ru3jt.jpg